1. Use the Midpoint Rule with n = 6 to approximate the integrale dx . Round the final answer to 6 decimal places.
Answer #1

2 - 6 mid pont Role : s tres de = or [ 4(**) *4803! * * $67)). 4 - I e da = I -25 – 49 – 81 - 121 e te te 69 -225 te 16 e To